OpenStudy (anonymous):

How would you write 0.3y + 1.2 = –0.4x in standard form?

OpenStudy (whpalmer4):

Standard form is Ax + By = C, with A and B not zero. Rearrange your equation so x and y are on one side and any leftover numbers are on the other.

OpenStudy (anonymous):

0.4x+0.3y=-1.2?

OpenStudy (whpalmer4):

Yes, I believe that is correct.

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

Standard form is handy if you want to quickly graph the line. at x = 0 (y-intercept), A(0) + By = C, so the y-intercept is just C/B. Similarly, for the x-intercept (y = 0) Ax = C, so the x-intercept is just C/A. Plot (0,C/B) and (C/A,0), draw a straight line through them, and you are done! For example: \[y = x + 2\]\[-1x + 1y = 2\] A = -1, B = 1, C = 2 Plot points (0,2) and (-2,0), draw line through them: